• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 10:05:33 on June 24, 2010

    Entry number 328110

    keyword=APEX experiment status

    The lead glass for the right arm is in the process of being put in the stack. The big cabling work will start when the stack for the left arm will. Radiation levels on one side of the septa are close to the limit where work can be performed, the other side still needs to cool off and should most likely be good tomorrow morning so I will start to get the work for the O-ring organized and have beam tomorrow.